OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Beluga Fun Tours: Age Ticket Price Under 3 years free 3 to 12 years $34 Over 12 years $45 On a certain day, the company took 103 people on whale watching trips. There were 64 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(64 - x)34 + (103 - 64) ⋅ 45 = C (103 + x)34 + (103 -(64 + x)) ⋅ 45 = C (103 - x)34 + (103 - 64) ⋅ 45 = C (64 + x)34 + (103 -(64 + x)) ⋅ 45 = C

OpenStudy (anonymous):

1st option

OpenStudy (anonymous):

You can get the number of adults by subtracting 103 by the number of children (64). Take that number and apply the adult rate. From there you have the variable x that you are trying to determine for kids under 3.